# Quantum transport in an ambipolar InSb nanowire quantum dot device

### v1

@@ Abstract
In modern integrated-semiconductor circuits, the inclusion of both $n$-type and $p$-type transistors is not merely a matter of preference but an essential requirement to meet the diverse and complex functionalities demanded by contemporary electronic devices and systems. Similarly, a quantum semiconductor circuit composed of both $n$-type and $p$-type units also holds significant potential as a versatile component for future spintronics devices and quantum computing circuits. For instance, single-electron and single-hole semiconductor quantum dot devices have emerged as promising candidates for qubit realization due to their controllable dual-level spin-up and spin-down states, as evidenced in numerous studies Loss and DiVincenzo (1998); Petta et al. (2005); Nowack et al. (2007); Brunner et al. (2009); Hu et al. (2012); Pribiag et al. (2013). However, integrating both $n$-type and $p$-type quantum dots within a single quantum circuit presents challenges. The difficulty lies in fabricating high-quality quantum devices capable of accommodating both carrier types since materialization and fabrication processes are often tailored specifically for one carrier type during doping or other work-function engineering procedures Zheng et al. (2004); Jin et al. (2004); Wang et al. (2011); Heinze et al. (2002); Giovannetti et al. (2008); Léonard and Talin (2011).
Meanwhile, the utilization of narrow band-gap materials for engineering ambipolar devices is feasible. III-V compound semiconductors with narrow band gaps, particularly InAs and InSb, offer promising platforms that can accommodate both $n$- and $p$-type polarities Doornbos and Passlack (2010); Lei et al. (2023). Due to their exceptional properties such as small effective carrier masses, high carrier mobilities, large $g$-factors, and pronounced spin-orbit interactions, InAs and InSb have also garnered significant attention as potential platforms for topological quantum computing Mourik et al. (2012); Deng et al. (2016). The realization of ambipolar InAs or InSb devices can be achieved through gate tuning, further highlighting their versatility and potential in future quantum technologies.
@@
The Nb-InSb nanowire quantum dot-Nb device investigated in this study [Fig. 1(a)] is fabricated from a zincblende InSb nanowire segment of an epitaxially grown InSb/InAs heterostructure nanowire. The nanowire is deposited onto a heavily doped Si/SiO2 substrate, which features a pre-fabricated Ti/Au (5/100 nm) layer at the back side serving as a global gate electrode. Ti/Nb/Al (3/80/5 nm) contacts with a spacing of approximately 100 nm are sputtered on both ends of the InSb nanowire segment. The 3-nm Ti is a sticking layer, while the 5-nm Al layer protects the Nb from oxidation. It should be noted that prior to sputtering the Ti/Nb/Al contacts, the nanowire undergoes treatment in an ammonia sulfide solution to eliminate the native oxide layer and passivate the fresh semiconductor surface. Subsequently, measurements are conducted on the device within a wet 3He/4He dilution refrigerator with a base temperature of 25 mK. For more comprehensive information regarding material growth, device fabrication, and measurement setup, please refer to Ref. Deng et al., 2014 and its associated references.
@@
Fig. 2: Coulomb oscillations in $n$-type and $p$-type regimes. (a) Linear-response conductance $G$ measured for the device shown in Fig. 1 at a positively-biased back gate voltage regime, with $V_{\rm sd}$ fixed at 10 $\mu$V. The conductance peaks are attributed to Coulomb oscillations in the electron-transport regime. (b) Same with (a), but measured at a more negative back gate range, with $V_{\rm sd}$ fixed at 200 $\mu$V. The conductance peaks here can be ascribed to Coulomb oscillations in the hole-transport regime. (c) Addition energy extracted from Coulomb oscillations in panel (a) as a function of the number of electrons in the quantum dot. An even-odd alternation behavior of the addition energy is observed. (d) Addition energy extracted for Coulomb oscillations in panel (b). No even-odd alternation is visible in the hole-transport regime.
@@
Fig. 3: Characterization of the electron-transport regime. (a) Charge stability diagrams in the electron-transport regime. Successive Coulomb diamond structures are clearly shown. (b) A close-up view of (a) in log-scale, to illustrate the superconductivity-induced conductance peaks. (c) Quantum level evolution in magnetic field, corresponding to Fig. 2(a). Here, the Zeeman effect induced level-drifting is evident. Blue numbers are a few extracted $g$-factors. At zero magnetic field, Kondo-effect induced conductance enhancement in the odd-occupied regimes can also be seen. (d) Quantum level evolution in the magnetic field with multiple Kondo-peaks, where single-solid-lines indicate spin-down quantum levels, double-lines indicate spin-up quantum levels, and dashed-lines are Kondo-peaks. Note that panel (c) comprises two individual measurements, stitched at $V_{\rm bg}=5.5$ V (indicated by the dashed line), to relieve the gate voltage hysteresis effect due to large-range sweep. To compensate the gate hysteresis, the high-$V_{\rm bg}$ part is offset to the left by 157 mV.
@@
For odd filling numbers, the corresponding $E_{\rm add}$ is primarily determined by the electron charging energy, denoted as $E_{c}^{n}$ (where the superscript n indicates the electron-transport regime). Conversely, for even-occupation quantum dots, the $E_{\rm add}$ comprises both the $E_{c}^{n}$ term and the quantum energy difference between adjacent orbital levels, $\Delta\epsilon$. Based on this analysis, we can extract the $E_{c}^{n}$ term to be in the range of 4.5$\sim$8 meV and the $\Delta\epsilon$ to be 0.7$\sim$5.8 meV.
In contrast, the hole-transport regime does not exhibit any visible even-odd alternation, indicating that $E_{\rm add}$ is predominantly influenced by the hole charging energy, $E_{c}^{h}$, in this regime. Since the effective mass of InSb heavy holes in the valence band is significantly larger than that of conductance electrons and light holes, it suggests that the quantum dot is predominantly populated by heavy holes due to the small energy quantization Vurgaftman et al. (2001).
@@
Note that, owing to the superconducting nature of the Nb-based contacts, a low-conductance strip is observed in the low-$V_{\rm sd}$ region, as evident in Fig. 3(b). This low-conductance strip arises from the proximity effect, which induces a superconducting gap in the InSb nanowire segments located beneath the Nb-based electrodes. For a more comprehensive understanding of measurements within the superconductor gap, readers are referred to a similar device reported in Ref. Deng et al., 2014. However, within the context of this study, the contacts are treated as normal leads due to the presence of a soft-induced superconductor gap with a considerable density of states within the gap itself.
@@
Fig. 4: Characterization of the hole-transport regime. (a) Charge stability diagrams in the hole-transport regime, where 30 regular Coulomb-diamond structures are shown. (b)-(d) Close-up view of the dashed rectangles in panel (a) with corresponding colors. (e) The same with (d) but measured at $B=4$ T. Lead-state induced conductance peaks can be identified in (b)-(e), which are indicated by black dashed-lines. The bottom schematics represent the quantum dot configurations at the corresponding positions indicated by circled numbers ①-⑦ in the transport diagrams. (f) Lead-state energy (relative to the Fermi level of the lead) as a function of $V_{\rm bg}$, extracted by tracing lead-state induced peaks. Note there could be lead-states in both the left and the right leads, but we only picture and analyze one side for simplicity without loss of generality.
@@
The device is then characterized in a perpendicularly applied magnetic field (relative to the substrate). As depicted in Fig. 3(c), notable shifts in the quantum dot levels of the $n$-type regime are observed. When subjected to a weak magnetic field, the evolution of these levels is primarily governed by the Zeeman effect. Specifically, quantum levels associated with up (down) spins tend to shift towards lower (higher) energy states. The magnitude of the Zeeman shift is determined by the formula $E_{\rm Z}=\frac{1}{2}g^{*}\mu_{B}B$, where $g^{*}$ represents the Lande $g$-factor, $\mu_{B}$ denotes the Bohr magneton, and $B$ is the strength of the magnetic field. In Fig. 3(c), we have marked several spins (indicated by black arrows) and subsequently extracted the corresponding $g$-factors. These extracted values are notably large and vary depending on the level, aligning with previous findings reported in Ref. Nilsson et al., 2009. However, as the magnetic field intensity increases, the evolution of these levels becomes more intricate due to the influence of the level interaction.
@@
Another noteworthy aspect observed in the magnetic field evolution is the emergence of the Kondo effect. When the filling number of the quantum dot is odd, Kondo correlations occur and give rise to a conductance enhancement, even within the Coulomb blockade region. For example, in Fig. 3(c), we have marked several Kondo peaks (using red arrows) at zero magnetic field. As the magnetic field increases, these Kondo peaks gradually disappear as the Zeeman effect lifts the spin-degeneracy. Nevertheless, it is conceivable that Kondo correlations may re-emerge at higher magnetic fields when spins of differing orbital levels align oppositely. This phenomenon is exemplified in Fig. 3(d), where four Kondo peaks (indicated by dashed lines) become apparent at finite magnetic fields (at higher $V_{\rm bg}$ values). These finite-field Kondo peaks, along with the resonant tunneling peaks and Kondo ridges observed at zero magnetic field, form a distinctive “honeycomb” pattern.
@@
We have also observed high conductance peaks outside each Coulomb diamond, e.g., the peak marked by ② in Fig. 4(b). Conductance peaks outside Coulomb diamonds have been reported, which can be attributed to the tunneling process through excited states of the quantum dot Franceschi et al. (2000); De Franceschi et al. (2001). However, the peaks outside Coulomb diamonds in Fig. 4(b) can not be ascribed to excited states because the small $\Delta\epsilon$ of the $p$-type regime does not match the energy difference between the ground state and the excited state. Instead, by tracking their positions relative to the Coulomb diamond edge, we attribute the out-diamond peaks to the electron density of states (DOS) in the leads, i.e., the lead-states or the reservoir-states Möttönen et al. (2010); Björk et al. (2004).
@@
As shown in the insets ①$\sim$ of Fig. 4, the actual leads are the InSb nanowire segments under the Nb-based contacts [Fig. 1(c)-(d)]. Within the nanowire, the continuum approximation holds true only for the longitudinal direction, rendering its DOS quasi-one-dimensional (Q1D) with highly nonuniform characteristics. In the tunneling regime of the quantum dot, the current is directly proportional to the DOS of the reservoir at a given energy. The differential conductance can be significantly suppressed even at the edge of Coulomb blockade diamonds when the quantum dot level is aligned with the valley of the Q1D DOS (①), while peaks in differential conductance emerge when the quantum dot level aligns with spikes in Q1D DOS (②).
@@
The Q1D lead-states can also be tuned by the back gate, although the lever arm is much smaller than the dot-level lever arm due to the screening effect of the metallic contacts. The lead-states are therefore trackable in the successive Coulomb diamonds. Conductance peaks originating from the same lead-state can emerge with different Coulomb diamonds, with a small energy drifting $\Delta E$ caused by $V_{\rm bg}$. The quantum dot can thus serve as a lead-state probe, and tomographically picture the dynamics of lead-states. The energy driftings for two lead-states are extracted and shown in Fig. 4(f). We can calculate the back gate lever arm for the lead-states to be about 0.005, which is about 5% of the lever arm for the dot states.
Moreover, we investigate the evolution of conductance peaks in response to a magnetic field. It is observed that the high conductance peak induced by lead states in Fig. 4(b) splits into two peaks in Fig. 4(e) as the magnetic field is increased to $B=4$ T. The schematic representation of this process can be seen in the inset sketches ⑥ and ⑦. Accordingly, we deduce the $g$-factor for the lead-state to be $g\sim$9.5. The extracted lead-state $g$-factor is significantly smaller than the dot-state $g$-factor in Fig. 3(c). The small $g$-factor can be ascribed to the superconductor screening effect from the Nb-based contacts. Even with an underestimation, the $g$-factor is still higher than that of InSb holes Pribiag et al. (2013). This is consistent with our assumption that the lead states are attributed to electron states in Nb-covered InSb nanowire segments.
@@
The Q1D lead-state picture in the $p$-type regime is also in line with the negative differential conductance observed in the $n$-type regime [Fig. 3(a)]. The sharp lead-state DOS profile can modifies the tunneling current significantly and gives rise to the negative differential conductance Möttönen et al. (2010).
@@
Acknowledgment. This work was financially supported by the Swedish Research Council (VR), the Ministry of Science and Technology of China (MOST) through the National Key Research and Development Program of China (No. 2016YFA0300601 and 2017YFA0303304), the National Natural Science Foundation of China (Nos. 11874071, 91221202 and 91421303), and the Beijing Academy of Quantum Information Sciences (No. Y18G22). M.T.D was supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302401).

### v2

@@ Abstract
Single-electron and single-hole semiconductor quantum dot devices have emerged as promising candidates for qubit realization due to their controllable two-level spin states (spin-up and spin-down), as evidenced in numerous studies [1, 2, 3, 4, 5, 6]. Transport measurements on electron and hole quantum dots provide valuable insights into the potential of these systems for implementing quantum computing. Realizing both single-electron and single-hole quantum dots within a single device through gating offers significant flexibility and innovation. However, integrating both $n$-type and $p$-type quantum dots into a single quantum circuit presents substantial challenges. The main difficulty lies in fabricating high-quality quantum devices that can accommodate both carrier types, as materialization and fabrication processes are often optimized specifically for one carrier type during doping or work-function engineering procedures [7, 8, 9, 10, 11, 12].
Meanwhile, the utilization of narrow band-gap materials for engineering ambipolar devices is feasible. III-V compound semiconductors with narrow band gaps, particularly InAs and InSb, offer promising platforms that can accommodate both $n$- and $p$-type polarities [13, 14]. Due to their exceptional properties such as small effective carrier masses, high carrier mobilities, large $g$-factors, and pronounced spin-orbit interactions, InAs and InSb have also garnered significant attention as potential platforms for topological quantum computing [15, 16, 17, 18, 19, 20, 21, 22]. The realization of ambipolar InAs or InSb devices can be achieved through gate tuning, further highlighting their versatility and potential in future quantum technologies.
@@
The Nb-InSb nanowire quantum dot-Nb device investigated in this study [Fig. 1(a)] has also been reported in the supplementary material of Ref. 18, where the superconductivity-related physics in the $n$-type region is discussed. To distinguish this device from the one described in the Appendix of the present article, we refer to the device in Fig. 1(a) as Device A. Device-A is fabricated from a zincblende InSb nanowire segment of an epitaxially grown InSb/InAs heterostructure nanowire. The nanowire is deposited onto a heavily doped Si/SiO2 substrate, which features a pre-fabricated Ti/Au (5/100 nm) layer at the back side serving as a global gate electrode. Ti/Nb/Al (3/80/5 nm) contacts with a spacing of approximately 100 nm are sputtered on both ends of the InSb nanowire segment. The 3-nm Ti is a sticking layer, while the 5-nm Al layer protects the Nb from oxidation. It should be noted that prior to sputtering the Ti/Nb/Al contacts, the nanowire undergoes treatment in an ammonia sulfide solution to eliminate the native oxide layer and passivate the fresh semiconductor surface. Subsequently, measurements are conducted on the device within a wet 3He/4He dilution refrigerator with a base temperature of 25 mK. For more comprehensive information regarding material growth, device fabrication, and measurement setup, please refer to Ref. 18 and its associated references.
@@
Fig. 2: Coulomb oscillations in $n$-type and $p$-type regimes. (a) Linear-response conductance $G$ measured for the device shown in Fig. 1 at a positively-biased back gate voltage regime, with $V_{\rm sd}$ fixed at 10 $\mu$V. The conductance peaks are attributed to Coulomb oscillations in the electron-transport regime. (b) Same with (a), but measured at a more negative back gate range, with $V_{\rm sd}$ fixed at 200 $\mu$V. The conductance peaks here can be ascribed to Coulomb oscillations in the hole-transport regime. The oscillation peak envelope (dashed-line) can be attributed to the shifting of the lead-states in the nanowire, see the main text below. (c) Addition energy extracted from Coulomb oscillations in panel (a) as a function of the number of electrons in the quantum dot. An even-odd alternation behavior of the addition energy is observed. (d) Addition energy extracted for Coulomb oscillations in panel (b). No even-odd alternation is visible in the hole-transport regime.
@@
Fig. 3: Characterization of the electron-transport regime. (a) Charge stability diagrams in the electron-transport regime. Successive Coulomb diamond structures are clearly shown. The spectrum from $V_{\rm bg}=4.5\sim 6.0$V is the same with Fig.5 in the supplementary material of Ref. 18. (b) A close-up view of (a) in log-scale, to illustrate the superconductivity-induced conductance peaks. (c) Quantum level evolution in magnetic field, corresponding to Fig. 2(a). Here, the Zeeman effect induced level-shifting is evident. Blue numbers are a few extracted $g$-factors. At zero magnetic field, Kondo-effect induced conductance enhancement in the odd-occupied regimes can also be seen. (d) Quantum level evolution in the magnetic field with multiple Kondo-peaks, where single-solid-lines indicate spin-down quantum levels, double-lines indicate spin-up quantum levels, and dashed-lines are Kondo-peaks. Note that panel (c) comprises two individual measurements, stitched at $V_{\rm bg}=5.5$ V (indicated by the dashed line), to relieve the gate voltage hysteresis effect due to large-range sweep. To compensate the gate hysteresis, the high-$V_{\rm bg}$ part is offset to the left by 157 mV.
@@
For odd filling numbers, the corresponding $E_{\rm add}$ is primarily determined by the electron charging energy, denoted as $E_{c}^{n}$ (where the superscript n indicates the electron-transport regime). Conversely, for even-occupation quantum dots, the $E_{\rm add}$ comprises both the $E_{c}^{n}$ term and the quantum energy difference between adjacent orbital levels, $\Delta\epsilon$. Based on this analysis, we can extract the $E_{c}^{n}$ term to be in the range of 2.3$\sim$6 meV and the $\Delta\epsilon$ to be 1.2$\sim$6.1 meV.
In contrast, the hole-transport regime does not exhibit any visible even-odd alternation, indicating that $E_{\rm add}$ is predominantly influenced by the hole charging energy, $E_{c}^{h}$, in this regime. We can extract $E_{c}^{h}\approx 4.2\sim 4.7$meV from Fig. 2(d). Since the effective mass of InSb heavy holes in the valence band is significantly larger than that of conductance electrons and light holes, it suggests that the quantum dot is predominantly populated by heavy holes due to the small energy quantization [23].
@@
Note that, owing to the superconducting nature of the Nb-based contacts, a low-conductance strip is observed in the low-$V_{\rm sd}$ region, as evident in Fig. 3(b). This low-conductance strip arises from the proximity effect, which induces a superconducting gap in the InSb nanowire segments located beneath the Nb-based electrodes. For a more comprehensive understanding of measurements within the superconductor gap for the device, see the supplementary material of Ref. 18.
@@
Fig. 4: Characterization of the hole-transport regime. (a) Charge stability diagrams in the hole-transport regime, where 30 regular Coulomb-diamond structures are shown. (b) and (c) Close-up view of the dashed rectangles in panel (a) with corresponding colors, where Lead-state induced conductance peaks can be identified (black dashed-lines). (d) Schematics represent the quantum dot configurations at the corresponding positions indicated by circles with colors in the transport diagrams. (e) Lead-state energy (relative to the Fermi level of the lead) as a function of $V_{\rm bg}$, extracted by tracing lead-state induced peaks. Note there could be lead-states in both the left and the right leads, but we only picture and analyze one side for simplicity without loss of generality.
@@
The device is then characterized in a perpendicularly applied magnetic field (relative to the substrate). As depicted in Fig. 3(c), notable shifts in the quantum dot levels of the $n$-type regime are observed. When subjected to a weak magnetic field, the evolution of these levels is primarily governed by the Zeeman effect. Specifically, quantum levels associated with up (down) spins tend to shift towards lower (higher) energy states. The magnitude of the Zeeman shift is determined by the formula $E_{\rm Z}=\frac{1}{2}g^{*}\mu_{B}B$, where $g^{*}$ represents the Lande $g$-factor, $\mu_{B}$ denotes the Bohr magneton, and $B$ is the strength of the magnetic field. In Fig. 3(c), we have marked several spins (indicated by black arrows) and subsequently extracted the corresponding $g$-factors. These extracted values are notably large and vary depending on the level, aligning with previous findings reported in Ref. 24. However, as the magnetic field intensity increases, the evolution of these levels becomes more intricate due to the influence of the level interaction.
Another observation in the evolution of the magnetic field is the appearance of high-conductance ridges when the filling number of the quantum dot is odd, which aligns with signatures of the Kondo effect [25, 26, 27, 28]. For instance, in Fig. 3(c), several high-conductance ridges are highlighted by red arrows at zero magnetic field. As the magnetic field increases, these ridges progressively vanish, consistent with the suppression of Kondo correlations due to the lifting of spin degeneracy by the Zeeman effect. Similar high-conductance ridges have been observed in other InAs and InSb nanowire quantum dot devices [29, 24, 30, 31, 18, 32]. Notably, in a comparable InSb quantum dot device reported in Ref. 32, temperature-dependent measurements confirmed that these high-conductance ridges are attributed to the Kondo effect.
@@
Nevertheless, it is conceivable that Kondo correlations may re-emerge at higher magnetic fields when two different orbital levels with opposite spins align. We also observed the re-emergence of the high-conductance ridges in Fig. 3(d), where four ridges (indicated by dashed lines) become apparent at finite magnetic fields (at higher $V_{\rm bg}$ values). The ridges, along with the resonant tunneling peaks, form a “honeycomb” pattern. The ridges are highly likely re-emerged Kondo peaks in the magnetic field [28].
@@
We have also observed high conductance peaks outside each Coulomb diamond, e.g., the peak marked by the red circle in Fig. 4(b). Conductance peaks outside Coulomb diamonds have been reported, which can be attributed to the tunneling process through excited states of the quantum dot [33]. However, the peaks outside Coulomb diamonds in Fig. 4(b) can not be ascribed to excited states because the small $\Delta\epsilon$ of the $p$-type regime does not match the energy difference between the ground state and the excited state. Instead, by tracking their positions relative to the Coulomb diamond edge, we attribute the out-diamond peaks to the electron density of states (DOS) in the leads, i.e., the lead-states or the reservoir-states [34, 35].
@@
As shown by the energy diagrams in Fig. 4(b), the actual leads are the InSb nanowire segments under the Nb-based contacts [Fig. 1(c)-(d)]. Within the nanowire, the continuum approximation holds true only for the longitudinal direction, rendering its DOS quasi-one-dimensional (Q1D) with highly nonuniform characteristics. In the tunneling regime of the quantum dot, the current is directly proportional to the DOS of the reservoir at a given energy. The differential conductance can be significantly suppressed even at the edge of Coulomb blockade diamonds when the quantum dot level is aligned with the valley of the Q1D DOS (labeled with yellow circle), while peaks in differential conductance emerge when the quantum dot level aligns with spikes in Q1D DOS (red circle).
@@
The Q1D lead-states can also be tuned by the back gate, although the lever arm is much smaller than the dot-level lever arm due to the screening effect of the metallic contacts. The lead-states are therefore trackable in the successive Coulomb diamonds. Conductance peaks originating from the same lead-state can emerge with different Coulomb diamonds, with a small energy shifting $\Delta E$ caused by $V_{\rm bg}$. The quantum dot can thus serve as a lead-state probe, and tomographically picture the dynamics of lead-states. The energy shiftings for two lead-states are extracted and shown in Fig. 4(e). The lead-state shifting also lead to an envelope for the Coulomb oscillation peaks in Fig. 2(b). We can calculate the back gate lever arm for the lead-states to be about 0.005, which is about 5% of the lever arm for the dot states.
@@
The Q1D lead-state picture in the $p$-type regime is also in line with the negative differential conductance observed in the $n$-type regime [Fig. 3(a)]. The sharp lead-state DOS profile can modifies the tunneling current significantly and gives rise to the negative differential conductance [34].
@@
Acknowledgment. This work was financially supported by the Swedish Research Council (VR), the Ministry of Science and Technology of China (MOST) through the National Key Research and Development Program of China (No. 2016YFA0300601 and 2017YFA0303304), the National Natural Science Foundation of China (Nos. 11874071, 91221202, 91421303, and 92165208), and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302401).
@@
Contribution. C.L.Y fabricated the devices. M.T.D and C.L.Y performed the measurements. P.C. grew the semiconductor nanowires. C.L.Y., M.T.D and H.Q.X analyzed the data. H.Q.X lead the experiments. All of the authors contributed to manuscript writing and discussion.
Appendix A Conductance spectra in magnetic fields
We have also investigated the evolution of lead-state induced conductance peaks in response to magnetic fields. As shown in Fig. 5, the high conductance peak at $B=0$T (labeled with the purple circle) seemingly splits into two peaks at $B=2$T, and further splits at $B=4$T. Although the peaks are weak at high magnetic fields, they can be cautiously attributed to the Zeeman splitting of the leads (cyan and yellow circles). Accordingly, we deduce the $g$-factor for the lead-state to be $g\sim$9.5. The extracted lead-state $g$-factor is significantly smaller than the dot-state $g$-factor in Fig. 3(c). The small $g$-factor can be ascribed to the superconductor screening effect from the Nb-based contacts. Even with an underestimation, the $g$-factor is still higher than that of InSb holes [6]. This is consistent with our assumption that the lead states are attributed to electron states in Nb-covered InSb nanowire segments.
Fig. 5: Charge stability diagrams in the $p$-type regime of device-A. (a-c) Charge stability diagram measured at $B=0$T, $B=2$T and $B=4$T, respectively. The lead-state induced high-conductance line (purple circle) splits into two weak lines at $B=2$T and further splits at $B=4$T (yellow and cyan circles). (d) Schematics of energy diagrams of the device.
Appendix B Data from device-B
Fig. 6: Ambipolar quantum transport measurements from another InSb quantum dot device (device-B). Main plot: Current measured in the linear response region as a function of back gate voltage. Coulomb oscillation peaks from both the $n$-type regime (the high peaks in the right side) and the $p$-type regime (the small peaks in the left side and the Coulomb stability in the inset) can be identified. The $p$-type Coulomb peaks start to show up when $V_{\rm bg}<-6$V and are several orders of magnitude lower than the $n$-type peaks. Charge-stability diagrams for the $p$-type regime is shown in the inset.
Signatures of ambipolar quantum transport are also observed from another InSb quantum dot device, denoted as device-B. Device-B has a geometry similar to that of device-A, but with Al leads. Device-B in superconducting state has been reported in Ref. 32.
As shown in Fig. 6, device-B also possesses an ambipolar quantum transport nature. As the gate voltage increases from around 0V to $-10$V, the device undergoes a transition from a $n$-type quantum dot to a $p$-type quantum dot. No Coulomb peaks are observed in the range of $V_{\rm bg}=-0.8\sim-6$V. Note that the $p$-type Coulomb peaks are several orders of magnitude lower than the $n$-type peaks, similar to device-A. Lead-states are difficult to identify for device-B due to the short lead length and the irregular Coulomb diamond relative to device-A.